10 years ago you placed $4000 into an account at Chase Bank. The account compounds interest continuously at a rate of 0.175% annually. How much money do you have in the account now?

Found 2 solutions by math_tutor2020, josgarithmetic:
Answer by math_tutor2020(3816)


Use the continuously compounded interest formula
A = P*e^(r*t)
where,
A = final amount
P = initial amount or deposit = 4000 dollars
e = special constant = 2.718...
r = annual interest rate = 0.00175
t = number of years = 10
Note: 0.175% = 0.175/100 = 0.00175

So,
A = P*e^(r*t)
A = 4000*e^(0.00175*10)
A = 4,070.61608860308
A = 4,070.62

Make sure to use your calculator's "e" button rather than using something like 2.718 to ensure the most accuracy possible. Your calculator will have more decimal digits stored to work with.
In this case, 4000*(2.718)^(0.00175*10) = 4,070.6087 approximately does round to 4,070.61; however, there may be other cases in which we aren't so lucky.

Answer: $4,070.62

Answer by josgarithmetic(39616)